# Classification and stability of penalized pinned elasticae

### v1

@@ 5. Elastic flow
(5.1) $\begin{cases}2\kappa^{\prime\prime}+\kappa^{3}-\lambda\kappa=0,\quad\text{in}\ \ (0,1),\\ \gamma(0)=(0,0),\ \gamma(1)=\gamma(\ell,0),\\ \kappa(0)=\kappa(1)=0.\end{cases}$
We first recall the following statement (see e.g. [21, Section 2] or [15, Proposition 5.2]):
@@
(5.2) $\omega$ :=$:=\Set{\gamma\in A_{\ell}}{\gamma\text{ is either $\gamma_{\rm seg}$, $\bar{\gamma}_{\rm larc}^{\lambda,\ell,1}$, or $\bar{\gamma}_{\rm sarc}^{\lambda,\ell,1}$}},$
@@
Since $L[\gamma_{\rm sarc}^{\lambda,\ell,1}]=2\mathrm{K}(q)/\alpha$, we see that $X^{\prime}(s)$ is increasing for $s\in(0,{L}/{2})$. Thus $X$ is convex in $[0,{L}/{2}]$. This together with the fact that $X(0)=0$ and $X(\frac{L}{2})=\ell/2$ implies that $X(s)<\ell/2$ holds for all $s\in(0,L/2)$. By reflection symmetry (cf. (2.21)), it suffices to check embeddedness of $\gamma_{\rm sarc}^{\lambda,\ell,1}(s)$ for $s\in(0,L/2)$, which immediately follows from monotonicity of $Y(s)=\frac{2q}{\alpha}\operatorname{cn}(\alpha s-\mathrm{K}(q))$ on $(0,L/2)$. ∎
@@
Let $\gamma(\cdot,t)$ be the smooth evolution of the $\lambda$-elastic flow in $A_{\ell}$ with initial datum $\gamma_{0}$. We prove the theorem by contradiction. If the asserted time $t_{0}$ does not exist, then one can find a sequence $t_{j}\rightarrow\infty$ such that $\gamma(\cdot,t_{j})$ is not embedded. By Proposition 5.1 we can find a subsequence $t_{j^{\prime}}\rightarrow\infty$ and reparametrizations $\phi_{j^{\prime}}:[0,1]\rightarrow[0,1]$ such that $\gamma(\phi_{j^{\prime}}(\cdot),t_{j^{\prime}})$ converges smoothly to some curve $\gamma_{\infty}$ satisfying (5.1). By (5.2), we can choose the reparametrizations $\phi_{j^{\prime}}$ in such a way that $\gamma_{\infty}$ must belong to $\omega$. In particular, $\gamma_{\infty}$ is embedded by Lemma 5.2. Since the set of embedded curves is open in the $C^{1}$-topology, (which can e.g. be seen like in [20, Lemma 4.3]) and $\gamma(\phi_{j^{\prime}}(\cdot),t_{j^{\prime}})\rightarrow\gamma_{\infty}$ in the $C^{1}$-topology, we obtain that $\gamma(\phi_{j^{\prime}}(\cdot),t_{j^{\prime}})$ must be embedded for some large $j^{\prime}$. Since reparametrizations do not affect the embeddedness of a curve, we also have that $\gamma(\cdot,t_{j^{\prime}})$ must be embedded. A contradiction. ∎
@@ Appendix A Elliptic integrals and functions
for $x\in\mathbf{R}$ and $q\in(0,1)$. Further we define $\mathrm{K}(q):=\mathrm{F}(\frac{\pi}{2},q)$ and $\mathrm{E}(q):=\mathrm{E}(\frac{\pi}{2},q)$. It is known that $[0,1)\ni q\mapsto\mathrm{K}(q)$ and $[0,1]\ni q\mapsto\mathrm{E}(q)$ are strictly increasing and strictly decreasing, respectively. More precisely, one has
@@
The function $Q:[0,1)\ni q\mapsto 2\mathrm{E}(q)-\mathrm{K}(q)$ is strictly decreasing, and satisfies $Q(0)=1$, $\lim_{q\to 1}Q(q)=-\infty$. In particular, there exists a unique $q_{*}\in(0,1)$ such that $02\mathrm{E}(q_{*})-\mathrm{K}(q_{*})=$ 0.
The constant $q_{*}$ stands for the modulus of the so-called figure-eight elastica; more precisely, the signed curvature of the figure-eight elastica is given by $k(s)=2q_{*}\operatorname{cn}(s,q_{*})$, up to similarity and reparametrization (cf. [20, Definition 5.1]).
@@
This follows from a straightforward calculation (cf. [25, Lemma 2.6]).
@@
Let $\operatorname{cn}(\cdot,q)$ and $\operatorname{sn}(\cdot,q)$ be the elliptic functions with modulus $q\in[0,1)$. Then, $\operatorname{cn}(\cdot,q)$ is an even $2\mathrm{K}(q)$-antiperiodic function on $\mathbf{R}$ and, in $[0,2\mathrm{K}(q)]$, strictly decreasing from $1$ to $-1$. Further, $\operatorname{sn}(\cdot,q)$ is an odd $2\mathrm{K}(q)$-antiperiodic function and in $[-\mathrm{K}(q),\mathrm{K}(q)]$ strictly increasing from $-1$ to $1$.
@@ Appendix B Technical proofs
Notice that by [20, Lemma B.4] the function $q\mapsto 2\frac{\mathrm{E}(q)}{\mathrm{K}(q)}-1$ is decreasing and positive on $(0,q^{*})$. The desired inequality follows then immediately from the previous two identities. Now we turn to the inequality $B[\gamma_{\rm sarc}^{\lambda,\ell,1}]>B[\gamma_{\rm larc}^{\lambda,\ell,1}]$. Proceeding as in (4.2) one can find
@@
we can rewrite $f(q)=a(q)\mathrm{K}(q)+b(q)\mathrm{E}(q)$. By Lemma A.1, $2\mathrm{E}(q)<\mathrm{K}(q)$ holds for $q\in[q_{*},1)$. This together with the fact that $a$ is negative on $[\tfrac{1}{\sqrt{2}},1]$ implies that, for $q\in[q_{*},1)$,
@@
Thus $f$ is strictly decreasing on $(\frac{1}{\sqrt{2}},q_{*})$, which yields uniqueness of $\hat{q}\in(\frac{1}{\sqrt{2}},q_{*})$ satisfying (2.4). Combining continuity of $f$ with (B.2) and (B.1), we see that $f>0$ on $[\frac{1}{\sqrt{2}},\hat{q})$ and $f<0$ on $(\hat{q},1)$. ∎
@@
Throughout this proof we write $m_{1}:=(\frac{3}{5})^{\frac{1}{2}}$ and $m_{2}:=(\frac{2}{3})^{\frac{1}{2}}$ for short and in the following we show $m_{1}<\hat{q}<m_{2}$. By definition of $f$ it follows that
@@
$=-\frac{14}{25}\int_{0}^{\frac{\pi}{2}}\Big{(}1-\frac{23}{14}+\frac{23}{14}\cdot\frac{3}{5}\sin^{2}\theta\Big{)}\frac{1}{\sqrt{1-\frac{3}{5}\sin^{2}\theta}}\;\mathrm{d}\theta$
$>-\frac{14}{25}\int_{0}^{\frac{\pi}{2}}\Big{(}-\frac{7}{14}+\sin^{2}\theta\Big{)}\frac{1}{\sqrt{1-\frac{3}{5}\sin^{2}\theta}}\;\mathrm{d}\theta>0,$
where in the last inequality we used the fact that, for any $q\in(0,1)$,
(B.4) $\begin{split}\int_{0}^{\frac{\pi}{2}}\frac{1-2\sin^{2}\theta}{\sqrt{1-q^{2}\sin^{2}\theta}}\;\mathrm{d}\theta&=\int_{0}^{\frac{\pi}{2}}\frac{\cos{2\theta}}{\sqrt{1-q^{2}\sin^{2}\theta}}\;\mathrm{d}\theta\\ &=\int_{0}^{\frac{\pi}{4}}\frac{\cos{2\theta}}{\sqrt{1-q^{2}\sin^{2}\theta}}\;\mathrm{d}\theta-\int_{0}^{\frac{\pi}{4}}\frac{\cos{2\theta}}{\sqrt{1-q^{2}\cos^{2}\theta}}\;\mathrm{d}\theta\\ &<0.\end{split}$
Thus we get $f(m_{1})>0$ and this together with Lemma 2.3 yields $m_{1}<\hat{q}$.
On the other hand, since (B.4) yields that
$\mathrm{E}(m_{2})-\frac{2}{3}\mathrm{K}(m_{2})=\frac{1}{3}\int_{0}^{\frac{\pi}{2}}\frac{1-2\sin^{2}\theta}{\sqrt{1-\frac{2}{3}\sin^{2}\theta}}\;\mathrm{d}\theta<0,$
we have $f(m_{2})=-\frac{5}{9}\mathrm{K}(m_{2})+\frac{7}{9}\mathrm{E}(m_{2})<-\frac{5}{9}\mathrm{K}(m_{2})+\frac{14}{27}\mathrm{K}(m_{2})<0$. This together with Lemma 2.3 gives $\hat{q}<m_{2}$ and the proof is complete. ∎
@@
where we used (1.2) and (A.1) in the last equality. For $q\in(\frac{1}{\sqrt{2}},\hat{q})$, we obtain (4.7) since $e^{\prime}=(\zeta^{\prime}\xi-\zeta\xi^{\prime})/\xi^{2}$. On the other hand, for $q\in(\hat{q},1)$, noting $e^{\prime}=-(\zeta^{\prime}\xi-\zeta\xi^{\prime})/\xi^{2}$ and the fact that $02\mathrm{E}(q)-\mathrm{K}(q)<$ 0, we see that (4.7) holds as well.

### v2

@@ 5. Elastic flow
(5.1) $\begin{cases}2\partial_{s}^{2}k+k^{3}-\lambda k=0\quad\text{in}\ \ (0,1),\\ \gamma(0)=(0,0),\ \gamma(1)=(\ell,0),\\ \kappa(0)=\kappa(1)=0.\end{cases}$
We first recall the following statement (see e.g. [NO_2014, Section 2] or [MPP21, Proposition 5.2]):
@@
[…]
@@
Since $L[\gamma_{\rm sarc}^{\lambda,\ell,1}]=2\mathrm{K}(q)/\alpha$, we see that $X^{\prime}(s)$ is increasing for $s\in(0,{L}/{2})$. Thus $X$ is convex in [0,L/2][0,{L}/{2}]. This together with the fact that $X(0)=0$ and $X(\frac{L}{2})=\ell/2$ implies that $X(s)<\ell/2$ holds for all $s\in(0,L/2)$. By reflection symmetry (cf. (2.21)), it suffices to check embeddedness of $\gamma_{\rm sarc}^{\lambda,\ell,1}(s)$ for $s\in(0,L/2)$, which immediately follows from monotonicity of $Y(s)=\frac{2q}{\alpha}\operatorname{cn}(\alpha s-\mathrm{K}(q))$ on $(0,L/2)$. ∎
@@
Let $\gamma(\cdot,t)$ be the smooth evolution of the $\lambda$-elastic flow in $A_{\ell}$ with initial datum $\gamma_{0}$. We prove the theorem by contradiction. If the asserted time $t_{0}$ does not exist, then one can find a sequence $t_{j}\rightarrow\infty$ such that $\gamma(\cdot,t_{j})$ is not embedded. By Proposition 5.1 we can find a subsequence $t_{j^{\prime}}\rightarrow\infty$ and reparametrizations $\phi_{j^{\prime}}:[0,1]\rightarrow[0,1]$ such that $\gamma(\phi_{j^{\prime}}(\cdot),t_{j^{\prime}})$ converges smoothly to some curve $\gamma_{\infty}$ satisfying (5.1). By (5.2), we can choose the reparametrizations $\phi_{j^{\prime}}$ in such a way that $\gamma_{\infty}$ must belong to $\omega$. In particular, $\gamma_{\infty}$ is embedded by Lemma 5.2. Since the set of embedded curves is open in the $C^{1}$-topology, (which can e.g. be seen like in [MR23, Lemma 4.3]) and $\gamma(\phi_{j^{\prime}}(\cdot),t_{j^{\prime}})\rightarrow\gamma_{\infty}$ in the $C^{1}$-topology, we obtain that $\gamma(\phi_{j^{\prime}}(\cdot),t_{j^{\prime}})$ must be embedded for some large $j^{\prime}$. Since reparametrizations do not affect the embeddedness of a curve, we also have that $\gamma(\cdot,t_{j^{\prime}})$ must be embedded. A contradiction. ∎
@@ Appendix A Elliptic integrals and functions
for $x\in\mathbf{R}$ and $q\in(0,1)$. Further we define $\mathrm{K}(q):=\mathrm{F}(\frac{\pi}{2},q)$ and $\mathrm{E}(q):=\mathrm{E}(\frac{\pi}{2},q)$. It is known that $0,1)\ni q\mapsto\mathrm{K}(q)$ and $0,1]\ni q\mapsto\mathrm{E}(q)$ are strictly increasing and strictly decreasing, respectively. More precisely, one has
@@
The function $Q:[0,1)\ni q\mapsto 2\mathrm{E}(q)-\mathrm{K}(q)$ is strictly decreasing, and satisfies $Q(0)=1$, $\lim_{q\to 1}Q(q)=$-$\infty$. In particular, there exists a unique $q_{*}\in(0,1)$ such that $2\mathrm{E}(q_{*})-\mathrm{K}(q_{*})=0$.
The constant $q_{*}$ stands for the modulus of the so-called figure-eight elastica; more precisely, the signed curvature of the figure-eight elastica is given by $k(s)=2q_{*}\operatorname{cn}(s,q_{*})$, up to similarity and reparametrization (cf. [MR23, Definition 5.1]).
@@
This follows from a straightforward calculation (cf. [Ydcds, Lemma 2.6]).
@@
Let $\operatorname{cn}(\cdot,q)$ and $\operatorname{sn}(\cdot,q)$ be the elliptic functions with modulus $q\in[0,1)$. Then, $\operatorname{cn}(\cdot,q)$ is an even $2\mathrm{K}(q)$-antiperiodic function on $\mathbf{R}$ and, in $0,2\mathrm{K}(q)]$, strictly decreasing from $1$ to 1-1. Further, $\operatorname{sn}(\cdot,q)$ is an odd $2\mathrm{K}(q)$-antiperiodic function and in $\mathrm{K}(q),\mathrm{K}(q)]$ strictly increasing from 1-1 to $1$.
@@ Appendix B Technical proofs
Notice that by [MR23, Lemma B.4] the function $q\mapsto 2\frac{\mathrm{E}(q)}{\mathrm{K}(q)}-1$ is decreasing and positive on $(0,q^{*})$. The desired inequality follows then immediately from the previous two identities. Now we turn to the inequality $B[\gamma_{\rm sarc}^{\lambda,\ell,1}]>B[\gamma_{\rm larc}^{\lambda,\ell,1}]$. Proceeding as in (4.2) one can find
@@
we can rewrite $f(q)=a(q)\mathrm{K}(q)+b(q)\mathrm{E}(q)$. By Lemma A.1, $2\mathrm{E}(q)<\mathrm{K}(q)$ holds for $q\in[q_{*},1)$. This together with the fact that $a$ is negative on $\tfrac{1}{\sqrt{2}},1]$ implies that, for $q\in[q_{*},1)$,
@@
[…]
@@
Throughout this proof we write $m_{1}:=(\frac{3}{5})^{\frac{1}{2}}$ and $m_{2}:=(\frac{2}{3})^{\frac{1}{2}}$ for short. First we show $m_{1}<\hat{q}$. In view of Lemma 2.3, it suffices to show $f(m_{1})>0$. To this end, here we adopt ideas from [Miura23, Proof of Lemma 3.10], to use the known series expansions of $\mathrm{K}$ and $\mathrm{E}$ (e.g. founded in [Abramowitz1992, 17.3.11, 17.3.2]) that are absolutely convergent
$\mathrm{K}(q)$ $=\int_{0}^{\frac{\pi}{2}}\frac{1}{\sqrt{1-q^{2}\sin^{2}\theta}}\;\mathrm{d}\theta=\frac{\pi}{2}\sum_{n=0}^{\infty}\bigg(\frac{(2n-1)!!}{(2n)!!}\bigg)^{2}q^{2n},$
$\mathrm{E}(q)$ $=\int_{0}^{\frac{\pi}{2}}\sqrt{1-q^{2}\sin^{2}\theta}\;\mathrm{d}\theta=\frac{\pi}{2}\sum_{n=0}^{\infty}\bigg(\frac{(2n-1)!!}{(2n)!!}\bigg)^{2}\frac{1}{1-2n}q^{2n}.$
We deduce from these expansions that
@@
$=\frac{9\pi}{50}\bigg[1-\sum_{n=1}^{\infty}A_{n}\Big(\frac{3}{5}\Big)^{n}\bigg],\quad\text{where}\quad A_{n}:=\bigg(\frac{(2n-1)!!}{(2n)!!}\bigg)^{2}\frac{28n+9}{9(2n-1)}.$
Since $A_{n}\leq A_{1}=\frac{37}{36}$ holds by induction, it follows that for all $N\in\mathbf{N}$
$\sum_{n=1}^{\infty}A_{n}\Big(\frac{3}{5}\Big)^{n}$ $\leq\sum_{n=1}^{N}A_{n}\Big(\frac{3}{5}\Big)^{n}+\sum_{n=N+1}^{\infty}\frac{37}{36}\Big(\frac{3}{5}\Big)^{n}$
$=\sum_{n=1}^{N}A_{n}\Big(\frac{3}{5}\Big)^{n}+\frac{185}{72}\Big(\frac{3}{5}\Big)^{N+1}=:T_{N}.$
An explicit computation (of finite operations multiplying integers) shows that
$T_{5}=\frac{570955201}{614400000}<1,$
and hence we obtain $f(m_{1})\geq\frac{9\pi}{50}(1-T_{5})>0$.
Now we prove $\hat{q}<m_{2}$ by showing that $f(m_{2})<0$. Noting that for any $q\in(0,1)$
(B.4) $\begin{split}\int_{0}^{\frac{\pi}{2}}\frac{1-2\sin^{2}\theta}{\sqrt{1-q^{2}\sin^{2}\theta}}\;\mathrm{d}\theta&=\int_{0}^{\frac{\pi}{2}}\frac{\cos{2\theta}}{\sqrt{1-q^{2}\sin^{2}\theta}}\;\mathrm{d}\theta\\ &=\int_{0}^{\frac{\pi}{4}}\frac{\cos{2\theta}}{\sqrt{1-q^{2}\sin^{2}\theta}}\;\mathrm{d}\theta-\int_{0}^{\frac{\pi}{4}}\frac{\cos{2\theta}}{\sqrt{1-q^{2}\cos^{2}\theta}}\;\mathrm{d}\theta\\ &<0,\end{split}$
we obtain
$\mathrm{E}(m_{2})-\frac{2}{3}\mathrm{K}(m_{2})=\frac{1}{3}\int_{0}^{\frac{\pi}{2}}\frac{1-2\sin^{2}\theta}{\sqrt{1-\frac{2}{3}\sin^{2}\theta}}\;\mathrm{d}\theta<0.$
This yields $f(m_{2})=-\frac{5}{9}\mathrm{K}(m_{2})+\frac{7}{9}\mathrm{E}(m_{2})<-\frac{5}{9}\mathrm{K}(m_{2})+\frac{14}{27}\mathrm{K}(m_{2})<0$. ∎
@@
[…]